# ToxicTags: Decoding Toxic Memes with Rich Tag Annotations

### v1

@@ 5 Analysis of the dataset
One of the most unique features of our dataset is the tags associated with each meme. This allows us to perform certain interesting analysis that we report below. Some of the most frequent tags we find in the train and the test splits are illustrated in Figure 1. We report the top 30 tags from the toxicity class as well as the top 10 tags each from the hateful, dangerous, offensive and normal classes, respectively. Some properties of the top tags are enlisted below (see Table 11 in the Appendix for more details).
@@
(ii) The most prevalent tags in the hateful class are ‘9/11’, ‘nazi’, ‘adolf hitler’, ‘pedophile’ and ‘racist’ indicating their popularity in spreading hateful sentiments. We also observe that for the dangerous class, some tags that appear benign at the surface level and are primarily associated with children’s media show up. These include ‘ernie and bert’, ‘sesame street’, and ‘sonic the hedgehog’ which are used to render a comic angle to the dangerous posts (see Table 10 in the Appendix for more examples).”
@@
6 Tag generation module
Figure 2: Schematic of the tag generation module with employed prompts and representative examples explaining each component.
A meme might not always be accompanied by tags, unlike in the case of our dataset. It is therefore important to have an automatic method to assign tags to an arbitrary input meme, which could enhance toxicity detection. Here, we propose a novel generative AI based architecture for obtaining socially relevant tags for a given input meme. Since directly obtaining tags from an input meme is not straightforward, we decompose the task into two steps. First, we generate an intermediate summary describing the input meme; next, we extract keywords from this summary, constituting the final set of tags. This design is inspired by the KeyLLM framework [^12] https://www.maartengrootendorst.com/blog/keyllm/. We next describe each component of the module in detail. Details on employed models and experimental setup are further discussed in Appendix F and H respectively.
Ground truth summary
We curate ground truth summaries that serve as an input for tag generation module discussed later. An overview of the ground truth generation setup is illustrated in Figure 2 (a). We prompt GPT-4o to generate the ground truth summary. We include the following items in the prompt so that a high-quality summary blended with the tags is generated.
(i) The meme’s metadata, e.g., the title and the OCR text.
(ii) The meme’s caption: to generate the image caption, we first apply LaMA image inpainting (Suvorov et al. 2021) to remove any OCR text from the image, and then prompt GPT-4o to produce a caption. This allows for a caption generation that is independent of the text embedded in meme.
(iii) The tags associated with the meme.
(iv) Enhanced context: VLMs being generative models may lack awareness of post-training developments, hence we supplement them with (a) a brief contextual description derived from Google Lens meme search results and (b) an expansion of each of the tags obtained using Google Search API. Details for collecting Google Lens context and its preprocessing are provied in Appendix C.
Manual verification: We randomly sampled 100 memes along with their corresponding GPT-4o generated summaries to conduct a manual verification. Three annotators judged each of the 100 samples where each of them was asked to rate the summaries on a scale of 1 to 10 in terms of the following three criteria: (i) completeness, (ii) fluency, and (iii) grammatical correctness. The evaluation yielded very high average scores of 7.89, 8.48 and 9.15 for completeness, fluency, and grammatical correctness, respectively, hence demonstrating that GPT-4o generated summaries can indeed serve as the ground truth for the pipeline.
Automatic summary generation
We now attempt to automatically generate summaries by fine-tuning various pretrained VLMs – PaliGemma-10B, LLaVA-1.6, and Qwen-2.5. For the fine-tuning, we use the training split noted in Table 2. The fine-tuning prompt queries the VLMs to generate a summary based on the input meme along with the metadata, the image caption and the Google Lens context. Note that we completely hide the tag information from the input. The generated summary is compared with the ground truth summary obtained in the previous section and the generation loss is back propagated. We use PEFT (Mangrulkar et al. 2022) with LoRA adapters (Hu et al. 2021) to support fine-tuning in a resource constrained scenario. The expectation is that this fine-tuning shall enforce the VLMs to learn to include the tags in the generated summary that are already blended in the ground truth summary.
Summary generation performance: We assess the generation performance using the test split noted in Table 2. For each test meme, we compare the generated summary with the corresponding ground truth summary using a suite of standard evaluation metrics, including Bleu score, chrF (character F1), Meteor, Sentence-BERT score, and Rouge. The results are noted in Table 3. We observe that across a majority of the metrics, PaliGemma-10B outperforms the other VLMs. Hence, for the purpose of the final tag extraction module in the next section, we shall use the summaries generated by the PaliGemma-10B model.
Model Bleu chrF MT S-Bert Rouge
PaliGemma-10B 0.1038 40.38 0.3188 77.65 0.2682
Qwen-2.5 0.0821 42.30 0.3156 76.67 0.2443
LLaVA-1.6 0.0292 18.10 0.1257 42.03 0.1117
Table 3: Performance of different VLMs in generating tagged summaries across different metrics. Best results are highlighted. Here, MT: Meteor and S-Bert: Sentence-BERT.
Tag generation
We fine-tune the different VLMs – PaliGemma-10B, LLaVA-1.6, and Qwen-2.5 for generating the final tags. The fine-tuning prompt asks these VLMs to extract a set of tags from the summary of the meme generated by PaliGemma-10B and fed as a part of the prompt. The extracted output tags are compared with the ground truth tags of the meme and the loss is back propagated. The fine-tuning is done again on the training split noted in Table 2. Similarly, we test the performance of the tag generation model on the test split noted in Table 2. In the following we first discuss the baselines and the metrics used for evaluation.
Baselines: We use three state-of-the-art baselines for evaluation. These include (i) Tag2Text (Huang et al. 2024), (ii) RAM (Zhang et al. 2024) and (iii) RAM++ (Huang et al. 2023). Tag2Text, is a vision-language pre-training framework that incorporates image tagging to enhance the learning of visual-linguistic representations. Tag2Text extracts tags directly from associated text, enabling the model to learn an image tagger while simultaneously guiding the vision-language learning process. The RAM model performs image tagging by training on large-scale image-text pairs. An initial model is trained by jointly learning from image captions and tags, supervised respectively by the original textual data and the parsed tags. This is followed by a data refinement engine that generates new tags and removes noisy annotations. The model on this cleaned dataset is retrained and further fine-tuned on a smaller, high-quality set. RAM++ is an open-set image tagging model that effectively harnesses multi-granular textual supervision. RAM++ unifies individual tag-level and global text-level supervision within a single alignment framework. To further enhance tag understanding, it leverages LLMs to transform semantically narrow tag supervision into broader, descriptive tag supervision, thereby enriching the model’s grasp of visual concepts in open-set scenarios. Both RAM and RAM++ models can either use their own trained tags or the open-set vocabulary of tags during inference, resulting in two different variants.
Metrics: We use multiple metrics to evaluate the similarity between the ground truth and the generated tags. These include the semantic similarity [^13] https://sbert.net/docs/sentence˙transformer/usage/semantic˙textual˙similarity.html, BERTScore and ConceptNet similarity [^14] https://conceptnet.io/. For each ground truth tag, we compute the semantic similarity (or BERT score or ConceptNet similarity) with all the generated tags and take the maximum value among these. Next, we average this maximum over all the ground truth tags. We average these averages over all the test data points. In addition, to capture better context, we also expand both the ground truth tag list as well as the generated tag list using the Google Search API and compare the expanded list using the same semantic similarity and BERTScore [^15] ConceptNet is primarily designed for obtaining one-word/short-phrase to one-word/short-phrase relationships. Hence, we have not used it to measure similarity between the expanded tag lists..
Model Tag similarity Exp tag similarity
$S$ BS ConceptNet $S$ BS
PaliGemma-10B 63.9 93.55 51.40 54.5 95.44
Qwen-2.5 59.1 92.65 45.23 48.2 94.77
LLaVA-1.6 56.7 87.89 43.85 45.0 92.88
RAM++ 40.6 90.36 19.43 24.9 92.48
RAM++ Openset 27.8 82.22 8.54 12.8 91.57
RAM 40.5 90.57 19.41 24.8 92.48
RAM Openset 28.3 72.19 7.07 12.2 91.62
Tag2Text 37.3 89.84 17.31 21.2 92.27
@@ 6 Tag generation module
Table 4: Performance of the tag generation module. Best results are highlighted. $S$: Semantic similarity, BS: BERTScore, Exp: Expanded.
@@
Ground truth trainers, hitler
@@
Our approach hitler, trainers, nazi, just do it 911, twin towers
RAM++ blue, man, shoe , smile building, city
RAM++ Openset
@@
RAM blue , man , shoe building, city
RAM Openset
@@
Table 5: Comparative examples of tags generated by different models. We provide more examples in Table 7 in the Appendix.
Generation performance: The results of tag generation are presented in Table 4. The top three rows (our scheme) by far outperforms the state-of-the-art baselines in terms all the evaluation metrics. Among our models, PaliGemma-10B has the best tag generation performance. State-of-the-art models fail to capture the meme’s social context and their implicit meaning. They focus primarily on the visible objects in the image (refer Table 5; also refer Table 7 in Appendix for more examples).
7 Toxic meme detection
ToxicTags– In this section we assess the performance of toxic meme detection on the test split for both stages: Stage I and Stage II. Results portray the effectiveness of the tags in few-shot exemplar selection as we demonstrate below (see Table 6). Note that the details of the employed VLMs, embedding generation models, prompt and experimental setup are covered in Appendix FG and H respectively.
Models Shots, Stage $_{\textrm{r}} I$$_{\textrm{im}}$ $_{\textrm{gt}} I$$_{\textrm{pt}}$ $_{\textrm{im}\oplus\textrm{gt}} I$$_{\textrm{im}\oplus\textrm{pt}}$
2, I 51.95 56.94 56.83 58.19 59.66 61.03
2, II 17.59 28.39 24.23 23.79 29.7 29.27
4, I 46.5 62.18 58.51 60.44 62.25 63
IDEFICS-3 4, II 17.58 39.27 35.52 35.45 41.6 41.4
2, I 64.4 66.89 68 67.45 68.56 69.12
2, II 44.29 50.68 53.28 51.41 55.6 54.47
4, I 65.97 70.17 68.86 70.44 70.54 70.34
GPT-4o 4, II 46.10 54.50 52.63 55.27 57.9 57.21
IntMeme I – 45.85 II – 31.6
Table 6: Macro-F1 scores for different few-shot example selection strategies evaluated on the ToxicTags dataset. Further, the final row shows the results for the latest and most competing baseline IntMeme (Hee and Lee 2025) evaluated on the ToxicTags dataset. Best: bold, second best: underline). Overall best result for each stage is highlighted.
Shot selection: In the following we describe the various methods for few-shot selection.
@@ 7 Toxic meme detection
[…]

### v2

@@ 5 Analysis of the dataset
One of the most unique features of our dataset is the tags associated with each meme. This allows us to perform certain interesting analysis that we report below.
Frequent tags: Some of the most frequent tags we find in the train and the test splits are illustrated in Figure 2. We report the top 30 tags from the toxicity class as well as the top 10 tags each from the hateful, dangerous, offensive and normal classes, respectively. The key observations are as follows.
@@
(ii) The most prevalent tags in the hateful class are ‘9/11’, ‘nazi’, ‘adolf hitler’, ‘pedophile’ and ‘racist’ indicating their popularity in spreading hateful sentiments. We also observe that for the dangerous class, some tags that appear benign at the surface level and are primarily associated with children’s media show up. These include ‘ernie and bert’, ‘sesame street’, and ‘sonic the hedgehog’ which are used to render a comic angle to the dangerous posts (see Table 12 in the Appendix for more examples)."
@@
Related tags: To understand the semantic themes associated with frequently occurring tag pairs, we conduct a co-occurrence analysis. Prior to the analysis, all tags are lemmatised to ensure consistency and improve result accuracy. Our primary objective is to identify how often specific tags co-occur and what thematic contexts they represent within the meme content. Some of the top co-occurring pairs and their themes are listed in Table 3.
Our study reveals that a large portion of frequent tag pairs relate to the Historical events/World War II theme, which indicates that memes are often coded around historical events for satire or shock value. From Table 3, we also observe that, to create humour or evoke strong reactions memes often engage with religious contrast/satire themes and controversial cultural commentary, such as 9/11 and Islamophobic references. Several pairs relate to themes such as nuclear warfare/World War II, child-related violence or abortion, creepy and disturbing content, and U.S. gun violence, indicating that memes frequently present serious or sensitive issues through humour or provocation. Themes such as Southern US stereotypes, taboo humour, humanitarian crises in Africa, and racism encode racial, regional, or humanitarian stereotypes, reflecting controversial social issues. Cartoon/anime character themes are often used in dark contexts.
6 The StemTox framework
6.1 Entropy-guided layer selection objective
Past approaches in toxic meme detection have focused primarily on single-task setups and rely either on zero-/few- shot prompting, or perform low-resource fine-tuning using stratetgies like PEFT/LoRA. Recently, multimodal research has evolved, and some works like He et al. (2025); Luo et al. (2024) propose a deeper understanding of the model’s internal representations and semantic alignment. A recent empirical evidence presented in Chen et al. (2025) suggests that deep representational alignment between vision and language modalities peaks at intermediate depths before undergoing representational drift in terminal layers. Taking cues from these works, we propose a novel token confidence based approach for simultaneous tag generation and toxic meme classification. Below, we detail our approach.
Figure 3 illustrates the complete workflow of StemTox. We consider a multimodal model $\mathcal{M}$ that takes a text prompt $\mathcal{T}$ and an image $\mathcal{I}$ as input and consists of $L$ stacked layers. Given an input sequence of fixed length, the model produces hidden representations across layers. Let $l\in\{1,\dots,L\}$ denote the layer index, $t\in\{1,\dots,T\}$ denote the token position in the input sequence, and $v\in V$ denote a token from the model vocabulary $V$. The hidden state $\mathbf{h}_{l,t}\in\mathbb{R}^{d_{h}}$ corresponds to token $t$ at layer $l$, as depicted in the figure, where $d_{h}$ represents the hidden dimension of the model. To process the model’s internal state, we extract token embeddings from all layers and project them into the model’s vocabulary space using the pretrained language model head $\mathbf{W}_{\text{vocab}}\in\mathbb{R}^{d_{h}\times|V|}$. This projection produces $\mathbf{w}_{l,t}$, which indicates the representation of the $t$-th token at layer $l$ in the vocabulary V:
$\mathbf{w}_{l,t}=\mathbf{h}_{l,t}\,\mathbf{W}_{\text{vocab}}.$ (1)
A softmax function is applied to $\mathbf{w}_{l,t}$ to obtain the confidence distribution $p_{l,t}$ over the vocabulary for the $t$-th token at layer $l$:
$p_{l,t}=\mathrm{softmax}\!\left(\mathbf{w}_{l,t}\right),$ (2)
To measure the confidence of the representation at layer $l$ for token $t$, we compute the Shannon entropy Vajapeyam (2014) $e_{l,t}$, which represents the entropy of token $t$ at layer $l$, as follows:
$e_{l,t}=-\sum_{v\in V}p_{l,t}(v)\,\log p_{l,t}(v).$ (3)
This entropy measure enables the dynamic selection of the least random token embedding across layers by selecting the representation corresponding to the minimum entropy, enabling the classification head to extract features from the most certain internal representations $\mathbf{h}_{\text{eff}}$:
$\mathbf{h}_{\text{eff}}=\mathrm{Aggregate}\!\left(\left\{\mathbf{h}_{l^{*},t}\;\middle|\;l^{*}=\arg\min_{l\in\mathcal{L}}e_{l,t}\right\}_{t=1}^{T}\right).$ (4)
Finally, the confident sequence representation $\mathbf{h}_{\text{eff}}$ is passed through the classification head to predict the class logits. The classification head consists of a sequence compression layer that compresses the sequence and projects it to a fixed-dimensional representation, which is then passed through an Multi-Layer Perceptron(MLP) for final classification:
$\textrm{classification label}=\mathrm{MLP}\!\left(\mathrm{Compress}\!\left(\mathbf{h}_{\text{eff}}\right)\right)$ (5)
We employ a weighted cross-entropy loss for the classification task, which is depicted as $\mathcal{L}_{\text{cls}}$ in the figure. To address class imbalance, class weights are computed following the weighting scheme described in King and Zeng (2001) [^12] https://scikit-learn.org/stable/modules/generated/sklearn.utils.class_weight.compute_class_weight.html.
approach model task/ shot toxicity detection tag similarity
stage I stage II chrF met $S$
mF1 acc mF1 acc
FS (TS) GPT 2-shot 61.26 75.37 48.59 51.70
4-shot 63.37 78.87 51.78 55.76
LHL PaliGemma uni 61.53 70.40 45.92 57.80
multi 70.48 82.20 61.54 69.20 0.4782 0.26 0.6220
LLaVA uni 61.58 74.60 44.69 58.70
multi 65.03 76.00 54.56 64.40 0.2585 0.1050 0.4540
StemTox PaliGemma uni 60.85 74.90 42.78 52.90
multi 72.55 83.90 64.66 72.40 0.4872 0.265 0.626
LLaVA uni 62.19 77.20 51.19 61.60
multi 67.24 77.70 57.99 66.90 0.2648 0.1070 0.4470
Table 4: Comparative results (reported out of 100) for toxicity detection and tag similarity across different approaches. The best results are highlighted. FS (TS): few-shot based on tag similarity; LHL: last hidden layer; mF1: Macro-F1; acc: Accuracy; chrF: chr-F; met: Meteor; $S$: Semantic-Similarity. (‘–’ indicates the task is not supported by methods.)
6.2 Multitask learning objective
A meme might not always be accompanied by tags, unlike in the case of our dataset. It is therefore important to have an automatic method to assign tags to an arbitrary input meme, which could enhance toxicity detection. As a combined objective, we therefore propose a novel multitasking framework where we jointly train toxic tag generation and discriminative label classification. The main motivation of this framework is to leverage semantic supervision to guide discriminative classification where the embeddings are shaped into a representation space where the supervision provided by the generative task acts as a regularizer for the shared representation space. Our experimental observations (refer Section 7) reveal that supervising the model to generate explicit toxic tags stimulates the intermediate hidden representations of the model and boosts discriminative performance, thereby signalling that the model learns more detailed and meaningful representations in joint training. The overall loss is calculated as a combination of the generation and classification losses:
$\mathcal{L}_{total}=\mathcal{L}_{gen}+\mathcal{L}_{cls}$ (6)
where, $\mathcal{L}_{gen}$ represents the generative loss and $\mathcal{L}_{cls}$ represents the classification loss. This alignment makes the layer selection objective described in Section 6.1 more robust as a toxicity classifier. The details on the employed models and experimental setup are elaborated in Appendix F and G respectively. In contrast, the uni-tasking framework (see Table 4), considers only the classification loss $\mathcal{L}_{\text{cls}}$, since it is solely dedicated for toxicity detection. In the next section we detail out our emperical observations.
7 Results
In this section we assess the performance of StemTox in toxic meme detection on the test split of the ToxicTags dataset for both stages – I and II. We also present the results of tag generation performance achieved by StemTox. For our experiments, we use two open-source VLMs – PaliGemma and LLaVA (refer Appendix F and G for further details). Table 4 shows the experimental results.
Toxic meme detection: StemTox with the multi-task learning objective consistently outperforms uni-task training across both models and stages, highlighting the benefit of joint learning for toxicity detection and tag generation, as summarized in Table 4. PaliGemma attains the highest macro F1 scores of 72.55 and 64.66 for stage I and II, respectively. This represents a notable improvement over the uni-task setting, which has macro F1 scores of 60.85 in stage I and 42.78 in stage II. A similar pattern is also observed for LLaVA, where StemTox in the multi-task framework achieves macro F1 scores of 67.24 and 57.99 for the two stages, outperforming the uni-task learning approch. Notably, StemTox outperforms the powerful GPT-4o model in a few-shot setup where the few-shot examples are selected based on the ground-truth tag similarity of the memes. To obtain these few-shot samples, we first compute the CLIP embeddings for each tag in the train set and also those that are associated with the query sample. Next, we find the maximum cosine similarity of each query tag embedding with the tag embeddings of an example meme from train set. For instance, given a test sample with tags $\{t_{1},t_{2},t_{3}\}$ and an example meme with tags $\{e_{1},e_{2},e_{3}\}$, we compute the cosine similarity between the CLIP embedding of tag $t_{1}$ and those of $e_{1}$, $e_{2}$, and $e_{3}$, and record the maximum similarity value. We repeat the same process for $t_{2}$ and $t_{3}$. The final tag similarity between the test and the example sample is then defined as the mean of the maximum similarity scores. Likewise StemTox also outperforms the setup where the last hidden layer is considered (LHL) for computing $\mathcal{L}_{cls}$ instead of the proposed entropy based scheme.
Competing baselines and generalizability across datasets: To further ground StemTox, we compare it with the existing benchmarks and frameworks. In particular, we compare with two benchmark datasets (FHM and MAMI) and five baselines namely, PromptHate Cao et al. (2023b), ProCap Cao et al. (2023a), ModHate Cao et al. (2024b), Few-Shot Hee et al. (2024) and M2KE Lu et al. (2025). Both the FHM and MAMI datasets do not have tags that is necessary to train StemTox. For the training points of both these datasets we therefore infer the tags using a PaliGemma model that is fine-tuned on the ToxicTags dataset to predict tags. Next we use the gold labels and these inferred silver tags to train StemTox (PaliGemma backbone as it exhibits the best performance in Table 4) for each of the two datasets. Table 5 compares the results of the baselines with StemTox. Our approach shows substantial improvement over the other baselines across the FHM and MAMI datasets, achieving a macro F1 score of 78.11, 79.67 and accuracy of 78.4, 80.1 respectively. In contrast, the best-performing baseline (M2KE) respectively achieves a maximum accuracy and macro F1 score of 75.76 and 75.52 on FHM, and 75.85 and 75.71 on MAMI.
Model FHM MAMI
acc mF1 acc mF1
PH 72.98 72.24 70.31 70.18
PC 75.10 74.85 73.63 73.42
MH 57.60 53.88 69.05 68.78
FS 66.00 65.80 70.50 70.10
MK 75.76 75.52 75.85 75.71
StemTox 78.4 78.11 80.1 79.67
Table 5: Results comparing our method with the previous works. Best results are highlighted. PH: PromptHate, PC: ProCap, MH: ModHate, FS: Few-Shot, MK: M2KE, mF1: Macro-F1, acc: Accuracy.
model tag similarity
chrF met $S$
RAM++ 0.1942 0.058 0.406
RAM++ (O) 0.094 0.012 0.279
RAM 0.19 0.05 0.40
RAM (P) 0.087 0.014 0.338
RAM (O) 0.079 0.013 0.285
Tag2Text 0.1631 0.041 0.372
Tag2Text (P) 0.2102 0.087 0.332
StemTox 0.4872 0.265 0.626
@@ 7 Results
Table 6: Performance of the tag generation module. Best results are highlighted. chrF: chr-F, met: Meteor, $S$: Semantic-Similarity, (O): open-set, (P): pre-trained using ToxicTags.
Tag generation module: As stated earlier, Table 4 presents the tag generation results over ToxicTags dataset as well. The metrics used are chr-F, Meteor and Semantic-Similarity [^13] https://huggingface.co/tasks/sentence-similarity. For each ground-truth tag, we compute each of the above metrics with all the generated tags and take the maximum value among these. Next, we average this maximum over all the ground-truth tags and calculate mean of these averages over all the test data points. Among the two models used within the StemTox multi-tasking framework, PaliGemma consistently outperforms LLaVA in terms of chr-F, Meteor and Semantic-Similarity. Specifically, PaliGemma achieves scores of 0.4872 (chr-F), 0.265 (Meteor) and 0.626 (Semantic-Similarity), whereas LLaVA attains 0.2648, 0.107 and 0.447, respectively. These results indicate that PaliGemma is more effective than LLaVA for tag generation in the multi-tasking framework StemTox.
Tag generation baselines: Since StemTox generate tags along with toxicity labels, it is imperative to compare its tag generation performance with state-of-the-art generators. To this purpose, we use three state-of-the-art baselines for evaluation namely Tag2Text Huang et al. (2024), RAM Zhang et al. (2024), and RAM++ Huang et al. (2023). To ensure fair evaluation, we further pre-trained the Tag2Text and RAM models on our ToxicTags dataset, allowing them to suitably adapt to domain-specific toxic vocabulary. Table 6 shows that StemTox by far outperforms all the other tag generation baselines in terms of all the three metrics – chr-F, Meteor and Semantic-Similarity. State-of-the-art models fail to capture the meme’s social context and their implicit meaning. They focus primarily on the visible objects in the image (refer to Table 7; also refer to Table 9 in Appendix for more examples).
@@
GT trainers, hitler
@@
ST hitler, trainers, nazi 911, jenga, twin towers
R++ blue, man, shoe , smile building, city
R++(O)
@@
R blue , man , shoe building, city
R(P)
R(O)
@@
T2T(P)
death, hitler, adolf hitler,
holocaust, family,
@@
Table 7: Examples of tags generated by different models. GT:Ground Truth, ST: StemTox, R: RAM, R++: RAM++, R(O): RAM open-set, R++(O): RAM++ open-set, T2T: Tag2Text, R(P): RAM pre-trained, T2T(P): Tag2Text pre-trained using ToxicTags. Please find more examples in Table 9 in the Appendix.
8 Error analysis
category PaliGemma LLaVA
hateful 0.2288 $\pm$ 0.2088 0.3029 $\pm$ 0.1968
dangerous 0.2423 $\pm$ 0.2043 0.3132 $\pm$ 0.1790
normal 0.3221 $\pm$ 0.1670 0.3478 $\pm$ 0.1557
offensive 0.4950 $\pm$ 0.1210 0.5559 $\pm$ 0.1065
Table 8: Average uncertainty scores ($\pm$ standard deviation) across different toxicity categories for PaliGemma and LLaVA. Higher numbers signify higher uncertainty in the model’s classification.
To understand the failures of the model we conduct an error analysis that helps to identify how these models struggle in classification; due to systematic weaknesses in contextual reasoning. We perform two complementary analyses to characterize these errors.
(i) We identify the most frequent tags for which each model struggles to predict the correct classification labels. Key observations: For LLaVA, the majority of the misclassifications are concentrated around tags corresponding to death-related events (death, funeral, suicide, murder, heart attack), mental health and violence (depression, gun control, slavery), and religious or cultural entities (religion, Jesus, Bible), indicating a bias toward emotionally sensitive scenes. In the case of PaliGemma, higher error frequencies are observed for tags associated with humor and internet slang (e.g., comic, lol, oof size large, cursed), suggesting that the model is unable to effectively interpret implicit sarcasm and meme culture.
(ii) To better understand where the models fail, we group all false positive samples according to their incorrectly predicted classes and perform an uncertainty-based error analysis. We quantify the model’s behavior using an uncertainty score derived from the Maximum Softmax Probability (MSP) Hendrycks and Gimpel (2017). This uncertainty score is defined as:
$\mathcal{U}(x)=1-\max_{c}p_{\theta}(c\mid x)$ (7)
where $x$ denotes a multimodal input sample, $c\in\{\textit{hateful},\textit{offensive},\textit{dangerous},\textit{normal}\}$ represents a toxicity class, and $p_{\theta}(c\mid x)$ is the model-predicted posterior probability given input $x$.
Key observations: Table 8 shows that LLaVA consistently exhibits higher uncertainty for false positive predictions. Conversely, PaliGemma shows lower confidence margins. Both models display the highest uncertainty in the offensive category, indicating difficulty in learning accurate patterns, whereas interestingly the lowest uncertainty is observed in the hateful category.